OpenStudy (anonymous):

Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent. Make sure you respond with complete sentences. Help !

OpenStudy (anonymous):

@wolfe8 can you help?

OpenStudy (anonymous):

@bibby can you?

OpenStudy (bibby):

Recall that you can write roots as fractional exponents like so:\[\huge \sqrt[3]{x^5} = x^{\frac{5}{3}}\]

OpenStudy (wolfe8):

First write out what are third root of x to the fourth power and I'm guessing x to the fifth

OpenStudy (anonymous):

that's the problem.. so would it be x4/3? im confused...

OpenStudy (wolfe8):

That's how your question is written? Then do as Bibby said to convert the root into a fractional power.

OpenStudy (anonymous):

x4/3...

OpenStudy (wolfe8):

A root of a number is the number to the power of the inverse of that root number.

OpenStudy (wolfe8):

Right. Then recall that when an exponent is outside a brackets you multiply the exponents.

OpenStudy (wolfe8):

Btw you should write that one as \[x ^{\frac{ 4 }{ 3}}\]

OpenStudy (anonymous):

so you'd multiply 4/3 & 5?

OpenStudy (wolfe8):

Yup

OpenStudy (anonymous):

\[20/3 ?\]

OpenStudy (anonymous):

@wolfe8

OpenStudy (wolfe8):

Yup

OpenStudy (anonymous):

so can you help me explain on how to find it.. in sentences :) please.

OpenStudy (wolfe8):

I did tell you what to do in sentences :)

OpenStudy (anonymous):

wanna put it all together for me? :)

OpenStudy (wolfe8):

>.> I think that's your part. Just put together what I said... convert the root into a fractional power. A root of a number is the number to the power of the inverse of that root number. when an exponent is outside a brackets you multiply the exponents.

OpenStudy (anonymous):

thank you (:

OpenStudy (wolfe8):

You're welcome. Good luck

OpenStudy (anonymous):

(: